OpenStudy (friendzone):

3 - 2(x - 4) = 5x + 3 While you are solving the equation, which choice could be one of your lines? A) 3x = 8 B) 3x = 14 C) 7x = 8 D) 7x = 14

OpenStudy (anonymous):

3-2x+8-5x-3=0 -7x+8=0 7x=8